# Effective Medium Theory for Time-modulated Subwavelength Resonators

### v1

@@ Abstract
This paper provides a general framework for deriving effective material properties of time-modulated systems of subwavelength resonators. It applies to subwavelength resonator systems with a general form of time-dependent parameters. We show that the resonators can be accurately described by a point-scattering formulation when the width of the resonators is small. In contrast to the static setting, where this point interaction approximation yields a Lippmann-Schwinger equation for the effective material properties, the mode coupling in the time-modulated case instead yields an infinite linear system of Lippmann-Schwinger-type equations. The effective equations can equivalently be written as a system of integro-differential equations. Moreover, we introduce a numerical scheme to approximately solve the system of coupled equations and illustrate the validity of the effective equation.
@@ 1 Introduction
This paper is devoted to deriving the homogenised equation of one-dimensional time-dependent metamaterials. The theory of homogenisation is central to a plethora of physical and engineering applications [17]. Homogenisation provides a way to transition from the detailed, small-scale description to an effective large-scale description, simplifying analysis and computation while preserving essential features. The homogenisation of a metamaterial with time-dependent parameters contributes significantly to the exploration of time-dependent materials. We specifically opt to study one-dimensional materials since they allow a more detailed exploration of the effect of time modulations. Several results have been shown in the one-dimensional time-dependent setting, such as the existence of space-time localised modes [4], the capacitance matrix approximation [1] and the rigorous definition of the scattered wave field based on a given incident wave field [2].
@@
Our theory in this paper provides a general framework for the effective medium theory of time-modulated subwavelength metamaterials. It applies to subwavelength resonator systems where the modulation frequency is of the same order as the subwavelength quasifrequencies and the operating frequency (i.e., the frequency of the incident wave) is in the low-frequency regime. It also generalises to higher space dimensions. In contrast to this paper, previous works have successfully established an effective medium theory for the low-frequency regime, in which the operating frequency is significantly smaller than the resonant frequency and the material parameters are static in time [6, 7, 15]. Furthermore, it is worth emphasising that the time-modulations considered in this paper are very different from the travelling wave-form modulations discussed in [16, 18, 14, 19] and in the references therein. By exploiting the subwavelength resonances of the system, they lead to fundamentally different wave propagation and scattering phenomena [2]. Moreover, their mathematical treatment is more involved. On the other hand, numerous papers have established an effective medium theory for subwavelength resonator systems based on the point interaction approximation [11, 10] resulting in the Lippmann-Schwinger equation [5, 9, 8, 3, 12]. In contrast to the static setting, we are faced with a coupled problem in the time-modulated case, which yields an infinite linear system of Lippmann-Schwinger-type equations instead. Moreover, implementing the point interaction approximation under the assumption of time-dependent material parameters is highly non-trivial, as there are two subwavelength quasifrequencies for a single time-modulated resonator. We shall see that, under suitable assumptions on volume fraction, configuration and incident frequency, we can adapt the results valid in the static case [5] to hold true in the time-modulated case.
@@
In this paper, we follow an approach involving the scattering matrix, which we shall define from first principles for both static and time-dependent metamaterials. We show that the scattering matrix has a characteristic structure in the leading-order terms as the resonators become small, which furnishes the point interaction approximation. We recall that the idea of a point interaction approximation goes back to Foldy’s paper [12]. It is a natural tool to analyse a variety of interesting problems in the continuum limit.
@@
Our point-scattering formulation is highly relevant in its own right, as the scattering matrix describes the coupling of different frequency harmonics and quantifies the frequency conversion due to the time-modulated parameters. Moreover, by taking the continuum limit, we derive the homogenised governing ordinary differential equation, which models the effective macroscopic behaviour while averaging out the fine-scale variations. This allows the introduction of a time-dependent one-dimensional effective medium theory, i.e., a theoretical framework to describe the macroscopic properties of heterogeneous time-dependent one-dimensional metamaterials in terms of their microscopic structure. The assumption of time-modulated material parameters in the derivation of an effective medium theory marks a new milestone in the mathematical exploration of metamaterials. Unlike the static case, we obtain a system of coupled Lippmann-Schwinger equations instead of a single equation [5]. This is due to the mode coupling that occurs as a consequence of time modulation.
@@
We start by providing an overview of the mathematical and physical setting of the problem considered herein in Section 2. In Section 3 we focus on one-dimensional metamaterials with static properties and derive the homogenised equation by introducing a scattering matrix and exploiting its structure. Although our main focus is on the time-modulated case, the static metamaterials serve as a concise introduction to our methods and ideas. This method is used to obtain an effective medium theory in Section 4 for time-modulated metamaterials. To achieve this, we explicitly compute the scattering matrix as the resonators become smaller. Finally, we solve the effective equation numerically in Section 5 with a uniquely tailored numerical scheme. We summarise our results and conclusions in Section 6.
@@ 2 Problem Setting
We consider a finite medium $\mathcal{U}\subset\mathbb{R}$, which contains $N$ disjoint high-contrast subwavelength resonators $\left(D_{i}\right)_{i=1,\dots,N}$. Each resonator is defined as an open interval $D_{i}:=\left(x_{i}^{-},x_{i}^{+}\right)$ of length $\ell_{i}:=x_{i}^{+}-x_{i}^{-}$ with a separation distance $\ell_{i(i+1)}:=x_{i+1}^{-}-x_{i}^{+}$, for all $i=1,\dots,N$. We denote the centre of each resonator by $z_{i}=(x_{i}^{-}+x_{i}^{+})/2$. In the remainder of this paper we shall assume that each resonator is of length $\ell$ and that they are evenly spaced. We refer to Figure 2.1 for an illustration of the geometrical setup of the material.
@@
In the acoustic setting we refer to $\rho$ as the material’s density and to $\kappa$ as the material’s bulk modulus. For our numerical simulations we shall consider
@@
for all $1\leq i\leq N$, where $\varepsilon_{\kappa,i}\in[0,1)$ are the amplitudes of the time-modulations and $\phi_{\kappa,i}\in[0,2\pi)$ the phase shifts.
@@
respectively. We consider a wave field $u(x,t)$ given by
@@
which is furnished by the assumption that $u(x,t)\mathrm{e}^{\mathrm{i}\omega t}$ is periodic with respect to $t$ with period $T$, where $\omega$ is the frequency of the incident wave. Note that the total wave field $u$ consists the scattered wave field and the incident wave field, in particular,
$u(x,t)=\begin{cases}u^{\mathrm{sc}}(x,t)+u^{\mathrm{in}}(x,t),&x\notin D,\\ u^{\mathrm{sc}}(x,t),&x\in D,\end{cases}$ (2.5)
and the modes of $u^{\mathrm{sc}}(x,t)$ and $u^{\mathrm{in}}(x,t)$ are given given by $v_{n}^{\mathrm{sc}}(x)$ and $v_{n}^{\mathrm{in}}(x)$, respectively. We shall give a closed form definition of the incident wave field in Section 3.
The governing equations posed on the modes $v_{n}$, can be derived to be the following [2]:
@@
$\left\{\begin{array}[]{ll}\displaystyle\frac{\mathrm{d}^{2}}{\mathrm{d}x^{2}}v^{\mathrm{sc}}_{n}+\frac{\rho_{0}(\omega+n\Omega)^{2}}{\kappa_{0}}v^{\mathrm{sc}}_{n}=0,&x\notin D,\\[10.00002pt] \displaystyle\frac{\mathrm{d}^{2}}{\mathrm{d}x^{2}}v_{n}^{\mathrm{sc}}+\frac{\rho_{\mathrm{r}}(\omega+n\Omega)^{2}}{\kappa_{\mathrm{r}}}v_{i,n}^{**}=0,&x\in D_{i},\quad i=1,\ldots,N,\\[10.00002pt] \displaystyle\left.v^{\mathrm{sc}}_{n}\right|_{-}\left(x_{i}^{-}\right)-\left.v^{\mathrm{sc}}_{n}\right|_{+}\left(x_{i}^{-}\right)=v_{n}^{\mathrm{in}}|_{-}(x_{i}^{-}),&\forall\,i=1,\dots,N,\\[5.0pt] \displaystyle\left.v^{\mathrm{sc}}_{n}\right|_{-}\left(x_{i}^{+}\right)-\left.v^{\mathrm{sc}}_{n}\right|_{+}\left(x_{i}^{+}\right)=-v_{n}^{\mathrm{in}}|_{+}(x_{i}^{+}),&\forall\,i=1,\dots,N,\\[5.0pt] \displaystyle\left.\frac{\mathrm{d}v_{n}^{\mathrm{sc}}}{\mathrm{~{}d}x}\right|_{+}\left(x_{i}^{-}\right)-\left.\delta\frac{\mathrm{d}v^{\mathrm{sc}}_{n}}{\mathrm{d}x}\right|_{-}\left(x_{i}^{-}\right)=\left.\delta\frac{\mathrm{d}v_{n}^{\mathrm{in}}}{\mathrm{d}x}\right|_{-}(x_{i}^{-}),&\forall\,i=1,\ldots,N,\\[10.00002pt] \displaystyle\left.\frac{\mathrm{d}v_{n}^{\mathrm{sc}}}{\mathrm{~{}d}x}\right|_{-}\left(x_{i}^{+}\right)-\left.\delta\frac{\mathrm{d}v^{\mathrm{sc}}_{n}}{\mathrm{d}x}\right|_{+}\left(x_{i}^{+}\right)=\left.\delta\frac{\mathrm{d}v_{n}^{\mathrm{in}}}{\mathrm{d}x}\right|_{+}(x_{i}^{+}),&\forall\,i=1,\ldots,N,\\[10.00002pt] \displaystyle\left(\frac{\mathrm{d}}{\mathrm{d}|x|}-\mathrm{i}\frac{(\omega+n\Omega)}{v_{0}}\right)v^{\mathrm{sc}}_{n}=0,&x\in(-\infty,x_{1}^{-})\cup(x_{N}^{+},+\infty).\end{array}\right.$ (2.6)
@@
We refer to [2, Appendix A] for a detailed derivation of (2.6). The wave number corresponding to the $n$-th mode inside and outside of the resonators is defined by
@@
respectively. The functions $v_{i,n}^{**}(x)$ are defined through the convolution
where $k_{i,m}$ are the Fourier series coefficients of $1/\kappa_{i}(t)$:
@@
The Fourier modes $v_{n}(x)$ of the wave field $u(x,t)$ solving (2.6) are given by
$v_{n}(x)=\begin{cases}\alpha_{n}^{i}\mathrm{e}^{\mathrm{i}k^{(n)}x}+\beta_{n}^{i}\mathrm{e}^{-\mathrm{i}k^{(n)}x},&\forall\,x\in\left(x_{i-1}^{+},x_{i}^{-}\right),\\ \sum\limits_{j=-\infty}^{\infty}\left(a_{j}^{i}\mathrm{e}^{\mathrm{i}\lambda_{j}^{i}x}+b_{j}^{i}\mathrm{e}^{-\mathrm{i}\lambda_{j}^{i}x}\right)f_{n}^{j,i},&\forall\,x\in\left(x_{i}^{-},x_{i}^{+}\right),\end{cases}$ (2.8)
where the coefficients $\alpha_{n}^{i},\,\beta_{n}^{i},\,a_{j}^{i},\,b_{j}^{i}$ need to be determined for all $i=1,\dots,N,\,j,n\in\mathbb{Z}$. The eigenpairs $\left(\lambda_{j}^{i},\mathbf{f}^{j,i}\right)_{i\in\mathbb{Z}}$ corresponding to the $i$-th resonator can be obtained as stated in [2, Lemma 2.1].
@@
The wave field $u(x,t)$ has infinitely many modes $v_{n}(x)$, which is problematic for the numerical implementation of the equations. Thus, we truncate the Fourier series by a truncation parameter $K\in\mathbb{N}$:
$u(x,t)=\sum\limits_{n=-K}^{K}v_{n}(x)\mathrm{e}^{-\mathrm{i}(\omega+n\Omega)t}.$ (2.9)

### v3

@@ Abstract
This paper provides a general framework for deriving effective material properties of one-dimensional, time-modulated systems of subwavelength resonators. It applies to subwavelength resonator systems with a general form of time-dependent parameters. We show that the resonators can be accurately described by a point-scattering formulation when the width of the resonators is small. In contrast to the static setting, where this point interaction approximation yields a Lippmann-Schwinger equation for the effective material properties, the mode coupling in the time-modulated case instead yields an infinite linear system of Lippmann-Schwinger-type equations. The effective equations can equivalently be written as a system of differential equations. Moreover, we introduce a numerical scheme to approximately solve the system of coupled equations and illustrate the validity of the effective equation.
@@ 1 Introduction
This paper is devoted to deriving the homogenised equation of one-dimensional, time-dependent metamaterials. The theory of homogenisation is central to a plethora of physical and engineering applications [23]. Homogenisation provides a way to transition from the detailed, small-scale description to an effective, large-scale description, simplifying analysis and computation while preserving essential features, and is a cornerstone method in the theory of metamaterials [11]. With significant attention devoted to metamaterials that are periodically driven in time, resulting in time-modulated material parameters, a natural problem is to develop homogenisation theories for time-dependent problems. We specifically choose to study one-dimensional materials since they allow a more detailed exploration of the effect of time-modulations. Several results have been established in the one-dimensional, time-dependent setting, notably the capacitance matrix approximation [2] and the rigorous definition of the scattered wave field for a given incident wave field [3].
This paper provides a general framework for the effective medium theory of time-modulated subwavelength metamaterials. The propagation of waves through these media is governed by the same equations as posed in [3]. Specifically, we assume that one of the material parameters $\kappa$ varies in time inside the resonators, while $\rho$ is a static material parameter (see Section 2 for a definition of the setting). We opt to only consider materials with time-dependent $\kappa$ and static $\rho$ since previous work has shown that modulating $\rho$ in time does not affect wave propagation at leading order [2]. Our theory applies to subwavelength resonator systems where the modulation frequency is of the same order as the subwavelength quasifrequencies and the operating frequency (i.e., the frequency of the incident wave) is in the low-frequency regime. It also generalises to higher space dimensions, as outlined in Remark 4.11.
@@
In contrast to this paper, previous works have successfully established an effective medium theory for the low-frequency regime, in which the operating frequency is significantly smaller than the resonant frequency and the material parameters are static in time [9, 10, 20]. Furthermore, it is worth emphasising that the time-modulations considered in this paper are very different from the travelling wave form modulations discussed in [22, 24, 19, 25] and in the references therein, where the mathematical analysis is simplified by passing to a moving coordinate frame. Opposed to the time-modulation in travelling wave form, our setting leads to the coupling of different frequency harmonics and is characterised by a system of coupled differential equations. By exploiting the subwavelength resonance of the building blocks, such systems may exhibit subwavelength resonant quasifrequencies [3], spatiotemporally localised modes [6], $k$-gaps and unidirectional band gaps [1]. Moreover, their mathematical treatment is more involved. In the static setting, numerous papers have established an effective medium theory for subwavelength resonator systems based on the point interaction approximation [15, 14] resulting in the Lippmann-Schwinger equation [7, 13, 12, 4, 16]. In contrast to the static setting, we are faced with a coupled problem in the time-modulated case, which yields an infinite linear system of Lippmann-Schwinger-type equations instead. Moreover, implementing the point interaction approximation under the assumption of time-dependent material parameters is highly non-trivial, as there are two subwavelength quasifrequencies for a single time-modulated resonator. We shall see that, under suitable assumptions on volume fraction, configuration and incident frequency, we can adapt the results valid in the static case [7] to hold true in the time-modulated case.
@@
In this paper, we follow an approach that involves the scattering matrix, which we shall define from first principles for both static and time-dependent metamaterials [21]. We show that the scattering matrix has a characteristic structure in the leading-order terms as the resonators become small, which furnishes the point interaction approximation. We recall that the idea of a point interaction approximation goes back to Foldy’s paper [16]. It is a natural tool to analyse a variety of interesting problems in the continuum limit. Taking the continuum limit allows us to derive a theory valid for infinitely many resonators, such as in [9] and [15], since it acts as a tool to average over an inhomogeneous material.
@@
Our point-scattering formulation is highly relevant in its own, as the scattering matrix describes the coupling of different frequency harmonics and quantifies the frequency conversion due to the time-modulated parameters. Moreover, by taking the continuum limit, we derive the homogenised governing ordinary differential equation, which models the effective macroscopic behaviour while averaging out the fine-scale variations. This allows the introduction of a time-dependent, one-dimensional effective medium theory, i.e., a theoretical framework to describe the macroscopic properties of heterogeneous time-dependent, one-dimensional metamaterials in terms of their microscopic structure. The assumption of time-modulated material parameters in the derivation of an effective medium theory marks a new milestone in the mathematical exploration of metamaterials, where classical wave-scattering results are generalised to the time-modulated setting. It is worth emphasising that while our purpose in [2, 3] was the study of wave scattering from systems of time-modulated systems of subwavelength resonators, our aim in the present paper is to derive effective models for computing the scattered fields by large systems of time-modulated resonators in the limit where the number of resonators goes to infinity and their typical size goes to zero, keeping the volume fraction of the resonators constant. Such non-classical models, which capture the average macroscopic behaviour of the large systems of resonators, can be used to simplify the analysis of time-modulated metamaterials and make the computations of their scattering properties more feasible and stable. Note also that because of the subwavelength resonant phenomena, the effective models both in the static case and in the time-modulated case depend on the operating frequency. Moreover, the effective models are not valid when the operating frequency coincides with the real part of one of the resonant frequencies [7]. In such situations, the effect of each resonator on the total scattered wave is of order one, and hence the limit does not exist as the number of resonators goes to infinity. The main difference between the static case and the time-modulated case comes from the fact that time-modulation gives rise to a family of coupled harmonics. This leads to an effective model described by a system of coupled differential equations instead of a Helmholtz equation with an effective (frequency dependent) potential as in the static case.
@@
Our paper is organised as follows. We start by providing an overview of the mathematical and physical setting of the problem considered herein in Section 2. In Section 3 we focus on one-dimensional metamaterials with static properties and derive the homogenised equation by introducing a scattering matrix and exploiting its structure. Although our main focus is on the time-modulated case, the static metamaterials serve as a concise introduction to our methods and ideas. This method is used to obtain an effective medium theory in Section 4 for time-modulated metamaterials. To achieve this, we explicitly compute the scattering matrix as the resonators become smaller. Finally, we solve the effective equation numerically in Section 5 with a uniquely tailored numerical scheme. We summarise our results and conclusions in Section 6.
@@ 2 Problem Setting
We consider a finite medium $\mathcal{U}\subset\mathbb{R}$, which contains $N$ disjoint high-contrast subwavelength resonators $\left(D_{i}\right)_{i=1,\dots,N}$. Each resonator is defined as an open interval $D_{i}:=\left(x_{i}^{-},x_{i}^{+}\right)$ of length $\ell_{i}:=x_{i}^{+}-x_{i}^{-}$ with a separation distance $\ell_{i(i+1)}:=x_{i+1}^{-}-x_{i}^{+}$, for all $i=1,\dots,N$. We denote the centre of each resonator by $z_{i}=(x_{i}^{-}+x_{i}^{+})/2$. In the remainder of this paper we shall assume that each resonator is of length $\ell$ and that they are evenly spaced. We denote the disjoint union of all resonators $D_{i}$ by $D$. We refer to Figure 2.1 for an illustration of the geometrical setup of the material.
@@
For our numerical simulations, we shall consider
@@
for all $1\leq i\leq N$, where $\Omega$ is the frequency of the time-modulations, $\varepsilon_{\kappa,i}\in[0,1)$ are their amplitudes, and $\phi_{\kappa,i}\in[0,2\pi)$ are the phase shifts.
@@
respectively.
We consider an incident wave field $u^{\mathrm{in}}(x,t)$ with a (real) frequency $\omega$ and such that
$\frac{\partial^{2}}{\partial t^{2}}u^{\mathrm{in}}(x,t)-v_{0}^{2}\Delta u^{\mathrm{in}}(x,t)=0.$ (2.4)
We let the total wave field $u(x,t)$ be given by
@@
which is furnished by the assumption that $u(x,t)\mathrm{e}^{\mathrm{i}\omega t}$ is periodic with respect to $t$ with period $T$, where $\omega$ is the frequency of the incident wave. We note that the total wave field $u$ consists of the scattered wave field and the incident wave field, in particular,
$u(x,t)=\begin{cases}u^{\mathrm{sc}}(x,t)+u^{\mathrm{in}}(x,t),&x\notin D,\\ u^{\mathrm{sc}}(x,t),&x\in D.\end{cases}$ (2.6)
Assuming that $u^{\mathrm{sc}}(x,t)\mathrm{e}^{\mathrm{i}\omega t}$ and $u^{\mathrm{in}}(x,t)\mathrm{e}^{\mathrm{i}\omega t}$ are periodic with respect to $t$ with period $T:=2\pi/\Omega$, we write their Fourier expansions with respect to $t$ as follows:
$u^{\mathrm{sc}}(x,t)=\sum\limits_{n=-\infty}^{\infty}v^{\mathrm{sc}}_{n}(x)\mathrm{e}^{-\mathrm{i}(\omega+n\Omega)t},\quad u^{\mathrm{in}}(x,t)=\sum\limits_{n=-\infty}^{\infty}v^{\mathrm{in}}_{n}(x)\mathrm{e}^{-\mathrm{i}(\omega+n\Omega)t},\quad\forall\,x\in\mathbb{R},\,t\geq 0.$ (2.7)
The governing equations for the wave scattering by the collection of resonators $D$ are given by [2, 5]:
@@
$\left\{\begin{array}[]{ll}\displaystyle\frac{\partial^{2}}{\partial t^{2}}u^{\mathrm{sc}}(x,t)-v_{0}^{2}\Delta u^{\mathrm{sc}}(x,t)=0,&x\notin D,\\[10.00002pt] \displaystyle\frac{\partial}{\partial t}\frac{1}{\kappa_{i}(t)}\frac{\partial}{\partial t}u^{\mathrm{sc}}(x,t)-v_{\mathrm{r}}^{2}\Delta u^{\mathrm{sc}}(x,t)=0,&x\in D_{i},\quad i=1,\ldots,N,\\[10.00002pt] \displaystyle u^{\mathrm{sc}}|_{-}(x_{i}^{-},t)-u^{\mathrm{sc}}|_{+}(x_{i}^{-},t)=u^{\mathrm{in}}|_{-}(x_{i}^{-},t),&i=1,\dots,N,\\[5.0pt] \displaystyle u^{\mathrm{sc}}|_{-}(x_{i}^{+},t)-u^{\mathrm{sc}}|_{+}(x_{i}^{+},t)=-u^{\mathrm{in}}|_{+}(x_{i}^{+},t),&i=1,\dots,N,\\[5.0pt] \displaystyle\frac{\partial u^{\mathrm{sc}}}{\partial x}\bigg{|}_{+}(x_{i}^{-},t)-\delta\frac{\partial u^{\mathrm{sc}}}{\partial x}\bigg{|}_{-}(x_{i}^{-},t)=\delta\frac{\partial u^{\mathrm{in}}}{\partial x}\bigg{|}_{-}(x_{i}^{-},t),&i=1,\dots,N,\\[10.00002pt] \displaystyle\frac{\partial u^{\mathrm{sc}}}{\partial x}\bigg{|}_{-}(x_{i}^{+},t)-\delta\frac{\partial u^{\mathrm{sc}}}{\partial x}\bigg{|}_{+}(x_{i}^{+},t)=\delta\frac{\partial u^{\mathrm{in}}}{\partial x}\bigg{|}_{+}(x_{i}^{+},t),&i=1,\dots,N,\\[10.00002pt] u^{\mathrm{sc}}\text{ is an outgoing wave,}\end{array}\right.$ (2.8)
where we use the notation
$\left.w\right|_{\pm}(x):=\lim_{s\rightarrow 0,\,s>0}w(x\pm s).$ (2.9)
We define the resonant quasifrequencies as the set of frequencies $\omega_{i}$ with real parts in the first Brillouin zone $[-\Omega/2,\Omega/2)$ such that there is a non-zero (Bloch) solution $u_{i}(x,t)$ to (2.8) with $u^{\textrm{in}}(x,t)=0$ and such that $u_{i}(x,t)\mathrm{e}^{\mathrm{i}\omega_{i}t}$ is periodic with respect to $t$ with period $T$. Such $\omega_{i}$ is a subwavelength resonant quasifrequency if the corresponding eigenmode is essentially supported in the subwavelength frequency regime, i.e., the infinite sum in (2.7) can be approximated by a finite one with a number of terms $K\ll 1/\sqrt{\delta}$, as $\delta\rightarrow 0$; see [2, 5].
A practical characterisation of these resonant quasifrequencies can be obtained by decomposing $u^{\mathrm{sc}}$ into its Fourier modes. In fact, the governing equations posed on the modes $v_{n}^{\mathrm{sc}}(x)$ and $v_{n}^{\mathrm{in}}(x)$ of $u^{\mathrm{sc}}(x,t)$ and $u^{\mathrm{in}}(x,t)$ defined in (2.7) can be derived to satisfy the following system of equations [3]:
$\left\{\begin{array}[]{ll}\displaystyle\frac{\mathrm{d}^{2}}{\mathrm{d}x^{2}}v^{\mathrm{sc}}_{n}+\frac{\rho_{0}(\omega+n\Omega)^{2}}{\kappa_{0}}v^{\mathrm{sc}}_{n}=0,&x\notin D,\\[10.00002pt] \displaystyle\frac{\mathrm{d}^{2}}{\mathrm{d}x^{2}}v_{n}^{\mathrm{sc}}+\frac{\rho_{\mathrm{r}}(\omega+n\Omega)^{2}}{\kappa_{\mathrm{r}}}v_{i,n}^{**}=0,&x\in D_{i},\quad i=1,\ldots,N,\\[10.00002pt] \displaystyle\left.v^{\mathrm{sc}}_{n}\right|_{-}\left(x_{i}^{-}\right)-\left.v^{\mathrm{sc}}_{n}\right|_{+}\left(x_{i}^{-}\right)=v_{n}^{\mathrm{in}}|_{-}(x_{i}^{-}),&\forall\,i=1,\dots,N,\\[5.0pt] \displaystyle\left.v^{\mathrm{sc}}_{n}\right|_{-}\left(x_{i}^{+}\right)-\left.v^{\mathrm{sc}}_{n}\right|_{+}\left(x_{i}^{+}\right)=-v_{n}^{\mathrm{in}}|_{+}(x_{i}^{+}),&\forall\,i=1,\dots,N,\\[5.0pt] \displaystyle\left.\frac{\mathrm{d}v_{n}^{\mathrm{sc}}}{\mathrm{~{}d}x}\right|_{+}\left(x_{i}^{-}\right)-\left.\delta\frac{\mathrm{d}v^{\mathrm{sc}}_{n}}{\mathrm{d}x}\right|_{-}\left(x_{i}^{-}\right)=\left.\delta\frac{\mathrm{d}v_{n}^{\mathrm{in}}}{\mathrm{d}x}\right|_{-}(x_{i}^{-}),&\forall\,i=1,\ldots,N,\\[10.00002pt] \displaystyle\left.\frac{\mathrm{d}v_{n}^{\mathrm{sc}}}{\mathrm{~{}d}x}\right|_{-}\left(x_{i}^{+}\right)-\left.\delta\frac{\mathrm{d}v^{\mathrm{sc}}_{n}}{\mathrm{d}x}\right|_{+}\left(x_{i}^{+}\right)=\left.\delta\frac{\mathrm{d}v_{n}^{\mathrm{in}}}{\mathrm{d}x}\right|_{+}(x_{i}^{+}),&\forall\,i=1,\ldots,N,\\[10.00002pt] \displaystyle\left(\frac{\mathrm{d}}{\mathrm{d}|x|}-\mathrm{i}\frac{(\omega+n\Omega)}{v_{0}}\right)v^{\mathrm{sc}}_{n}=0,&x\in(-\infty,x_{1}^{-})\cup(x_{N}^{+},+\infty),\end{array}\right.$ (2.10)
where the functions $v_{i,n}^{**}(x)$ are defined through the convolution
with $k_{i,m}$ being the Fourier series coefficients of $1/\kappa_{i}(t)$:
@@
We refer to [3, Appendix A] for a detailed derivation of (2.10). In the following, we denote the wave numbers corresponding to the $n$-th mode inside and outside of the resonators by
@@
respectively.
Remark 2.1.
While the incident frequency $\omega$ is real, the resonant frequencies $\omega_{i}$ of (2.10) are, in general, complex. Note that, as shown in [17], there may be real resonances. At such frequencies (known in physics as lasing points [18]), the scattering problem (2.10) is ill-posed and we do not expect solutions to be unique. For the remainder of this paper, we assume that all the resonant frequencies $\omega_{i}$ are strictly non-real, so that there exists a unique solution to (2.10) for real-valued $\omega$. Note that under this assumption, the well-posedness of (2.10) for $\omega$ real and of (2.8) are equivalent.
@@
The Fourier modes $v^{\mathrm{sc}}_{n}(x)$ of the wave field $u^{\mathrm{sc}}(x,t)$ solving (2.10) are given by
$v^{\mathrm{sc}}_{n}(x)=\begin{cases}\alpha_{n}^{i}\mathrm{e}^{\mathrm{i}k^{(n)}x}+\beta_{n}^{i}\mathrm{e}^{-\mathrm{i}k^{(n)}x},&\forall\,x\in\left(x_{i-1}^{+},x_{i}^{-}\right),\\ \sum\limits_{j=-\infty}^{\infty}\left(a_{j}^{i}\mathrm{e}^{\mathrm{i}\lambda_{j}^{i}x}+b_{j}^{i}\mathrm{e}^{-\mathrm{i}\lambda_{j}^{i}x}\right)f_{n}^{j,i},&\forall\,x\in\left(x_{i}^{-},x_{i}^{+}\right),\end{cases}$ (2.12)
where the coefficients $\alpha_{n}^{i},\,\beta_{n}^{i},\,a_{j}^{i},\,b_{j}^{i}$ need to be determined for all $i=1,\dots,N,\,j,n\in\mathbb{Z}$. The eigenpairs $\left(\lambda_{j}^{i},\boldsymbol{f}^{j,i}\right)_{i\in\mathbb{Z}}$ corresponding to the $i$-th resonator can be obtained as stated in [3, Lemma 2.1].
@@
The wave field $u(x,t)$ has infinitely many modes $v_{n}(x)$, which is problematic for the numerical implementation of the equations. Thus, we approximate $u$ by a truncated Fourier series with $K\in\mathbb{N}$ [2]:
$u(x,t)\approx\sum\limits_{n=-K}^{K}v_{n}(x)\mathrm{e}^{-\mathrm{i}(\omega+n\Omega)t}.$ (2.13)
Throughout the remainder of this work, we consider a truncation of (2.10) where we restrict to $-K\leq n\leq K$ and similarly seek a truncated $u$ of the form (2.13). This allows for numerical results, since otherwise, we would be dealing with an infinite coupled system. We note that the convergence theory and an estimation of the error introduced by truncating the Fourier series of $u$ remain open problems.